2. In a truck-loading station at a post office, a small 0.200-kg package is released from rest at

point A on a track that is one quarter of a circle with radius 1.60 m.

The size of the package is much less than 1.60 m, so the package can be treated as a particle. It

slides down the track and reaches point B with a speed of 4.80 m/s.

From point B, it slides on a level surface a distance of 3.00 m to point C, where it comes to rest.

(a) What is the coefficient of kinetic friction on the horizontal surface?

(b) How much work is done on the package by friction as it slides down the circular arc from A

to B?


1
Expert's answer
2021-10-10T16:00:01-0400

1)


"\mu mgL=0.5mv^2\\ (9.8)(3)\mu =0.5(4.8)^2\\\mu=0.39"

2)


"W=0.5(0.2)(4.8)^2-0.2(9.8)(1.6)=-0.83\ J"